# Flux-Activated Resonant Control of a Bosonic Quantum Memory

### v1

@@ Abstract
Universal control of bosonic degrees of freedom provides a hardware-efficient route for quantum information processing with high-dimensional systems. Bosonic circuit quantum electrodynamics (cQED), which leverages transmon ancillae to coherently control long-lived superconducting cavities, is well suited to this goal. However, the cavity transitions are nearly degenerate in the usual dispersive regime, which limits the direct addressability of individual excitation levels and increases the complexity of engineered gates. Here, we integrate an on-chip flux-control architecture with a long-lived bosonic memory housed in a 3D superconducting cavity to dynamically access resonant Jaynes–Cummings (JC) interactions, enabling universal control of the cavity field with an analytically simple protocol. We deterministically prepare Fock states and their superpositions by sequentially climbing the JC ladder, and realize efficient arbitrary rotations between any pair of Fock levels in the memory. This on-demand access to JC interactions offers a versatile toolbox for implementing robust Fock-basis qudits and harnessing the rich dynamics of high-dimensional bosonic elements for quantum information processing.
Effective control over a large Hilbert space is essential for both practical quantum information processing [47, 61, 18] and explorations of the rich underlying concepts of quantum physics [2, 17, 1]. The pursuit of progressively larger systems often takes the form of assembling collections of many interconnected and individually controlled two-level elements. This scaling incurs complex hardware overhead and imposes a significant challenge to current quantum technologies. One strategy to reduce hardware requirements is to use quantum systems that are intrinsically more complex, such as $d$-level systems (qudits) [30, 55]. This pathway has been explored in several platforms such as in molecular spins [44, 39], optical photons [56, 22], trapped ions [34, 46], silicon nuclear spins [60] and superconducting artificial atoms [57, 33, 8].
An alternative approach is to leverage the large Hilbert space of quantum harmonic oscillators. Notably, bosonic circuit QED systems use 3D superconducting cavities coupled to transmon ancillae to realize highly controllable and coherent oscillators. Such devices have demonstrated Fock state preparation up to 100 photons [10] and cat states with an average photon population of $1024$ [38].
@@
Universal control of the oscillator is typically achieved via dispersive coupling to a transmon qubit [4]. In this regime, cavity transitions are close in frequency (Fig. 1a) and not individually addressable. Therefore, arbitrary operations on the oscillator state often require driving the transmon and the cavity simultaneously with complex numerically optimized waveforms [26, 15, 14]. However, these numerical pulses are susceptible to parasitic dynamics, and their ‘black-box’ nature often complicates effective troubleshooting. Although control can also be implemented using higher-order parametric processes without numerical optimization [19, 23], they are limited by the achievable coupling strength or drive-induced errors [53, 36]. Developing analytical, versatile control methods that directly address individual cavity transitions would therefore fill an important gap for bosonic quantum information processing.
Figure 1: Comparison of the dispersive and resonant regimes. (a) The bosonic transitions in the dispersive regime are degenerate within the each qubit manifold and are separated by $\chi\sim$ MHz between the $|g\rangle$ and $|e\rangle$ manifold. (b) This degeneracy is lifted in the resonant JC regime, with level splittings $\propto g$$\sim$ tens of MHz, enabling fast selective excitation of individual bosonic levels. c) The frequency spectra of both coupling regimes shown on the same axis.
@@
In this work, we use flux-activated resonant Jaynes-Cummings (JC) interactions to access number-selective cavity transitions and implement arbitrary yet analytically simple control of a bosonic memory mode. These dynamics are enabled by a dynamically adjusted transmon-cavity detuning, which can shift the coupled system between a highly nonlinear resonant regime with well-resolved JC transitions and the dispersive regime with suppressed cavity photon loss. We protect the coherence of the bosonic memory against flux line leakage by incorporating on-chip low-pass filtering over the frequency band occupied by the cavity and the transmon. As a result, the lifetime of the bosonic mode reaches $T_{1}=590\pm 40\,\mu\textnormal{s}$ ($Q=(2.6\pm 0.2)\times 10^{7}$) when the transmon is tuned to be dispersively coupled – more than a two-fold improvement over previous implementations using the flux hose [52, 3]. We leverage these new capabilities to demonstrate the direct creation of Fock states and their superpositions by sequentially climbing the JC ladder (Fig. 1b). Furthermore, we achieve state transfer between any two chosen levels of the cavity by using multi-tone pulses and realize arbitrary rotations between them via simple analytical sequences. These results form the foundation for universal Fock-basis qudit control and robust quantum information processing using high-dimensional systems.
Figure 2: Architecture for robust flux-delivery in a high-Q bosonic circuit QED device. (a) High-Q $\lambda/4$ coaxial cavity coupled to an asymmetric SQUID-transmon (shown in bottom-left inset). The flux bias is delivered by a transmission line wirebonded to an SMA port at the opposite end of the chip. The flux line and a readout resonator partially adopt a coplanar waveguide (CPW) architecture. They are equipped with two 33rd-order Chebyshev low-pass filters in series (schematics shown in top-right inset) and a Purcell filter, respectively, to protect the cavity and the transmon against decay. (b) The cavity $T_{1}$ is measured while the transmon is parked at a few selected flux points with different values of $\Delta=\omega_{cav}-\omega_{T}$, showing an increase of memory lifetime $T_{1}$ as the transmon-induced decay is mitigated at larger detunings.
Our method to achieve universal control of the bosonic mode is based on adiabatically connecting the eigenstates of the resonant JC Hamiltonian to the product eigenbasis of the decoupled circuits [49]. The resonant JC Hamiltonian in the interaction picture is given by
$\hat{H}_{JC}=g\left(\hat{a}^{\dagger}\hat{\sigma}^{-}+\hat{a}\hat{\sigma}^{+}\right),$ (1)
@@
where $g=12.2\pm 0.1\,\textnormal{MHz}$ is the coupling factor, $\hat{a}$ ($\hat{a}^{\dagger}$) is the annihilation (creation) operator of the cavity and $\hat{\sigma}^{-}$ ($\hat{\sigma}^{+}$) is the lowering (raising) Pauli operators of the transmon taken as a two-level system. The $\hat{H}_{JC}$ eigenstates are the ground state $|0\rangle$ and the $N$th-excited levels
@@
$|N\pm\rangle=\frac{1}{\sqrt{2}}\left(|N,g\rangle\pm|N\text{-}1,e\rangle\right),$ (2)
@@
where $|N,s\rangle$ is the product of the cavity Fock state $|N\rangle$ and the transmon state $|s=g,e\rangle$, with eigenfrequencies $\omega_{|N\pm\rangle}=\pm\sqrt{N}g$. This photon-number dependence of $\omega_{|N\pm\rangle}$ leads to the direct addressability of the transitions as illustrated in Fig. 1b.
@@
$|N+\rangle\,\overset{\omega_{N,r}}{\longleftrightarrow}\,|(N\text{+}1)-\rangle,$ (3)
$|N-\rangle\,\overset{\omega_{N,b}}{\longleftrightarrow}\,|(N\text{+}1)+\rangle,$
@@
with transition frequencies $\omega_{N,r}=-\left(\sqrt{N+1}+\sqrt{N}\right)g$ and $\omega_{N,b}=\left(\sqrt{N+1}+\sqrt{N}\right)g$ relative to the rotating frame, respectively. This set of transitions, together with the ground state excitations $|0\rangle\leftrightarrow|1\pm\rangle$, can connect any two eigenstates of the resonant JC Hamiltonian with a high frequency selectivity scaling with $g$ (Fig.1c). When the transmon is adiabatically tuned out of resonance with the cavity to lower frequencies, the eigenstates in Eq. (2) transform into products of cavity and transmon eigenstates: $|N+\rangle\to|N,g\rangle$ and $|N-\rangle\to|N\text{-}1,e\rangle$. The first sideband in Eq. (3) is transformed into the transition $|N,g\rangle\leftrightarrow|N,e\rangle$, which is available in standard dispersive regime [45]. In contrast, the blue sideband, $|N\text{-}1,e\rangle\leftrightarrow|N\text{+}1,g\rangle$, is only directly (using a single drive) accessible with the resonant JC dynamics [49]. Therefore, combining direct access to the resonant JC dynamics and the ability to adiabatically detune the qubit from the oscillator, we can orchestrate any desired bosonic mode.
@@
To leverage this resonant control scheme, we construct a new bosonic circuit QED hardware that can simultaneously provide strong and broadband flux control while protecting the cavity from decay (Fig. 2a). Our device features a 3D superconducting cavity as the long-lived quantum memory, together with a lithographically defined on-chip flux line placed in close proximity to the SQUID loop with mutual inductance of $0.387\,\Phi_{0}/\textnormal{mA}$ with the SQUID loop ($470\,\mu\textnormal{m}^{2}$) and an average magnetic field of $1.70\,\mu\textnormal{T/mA}$. This highly effective flux delivery structure allows to minimize the loop area and hence reduce the impact of flux-noise on the transmon coherence times [50]. In this particular device, we measure transmon $T_{1}\approx 23\,\mu\textnormal{s}$ and $T_{2}^{\textnormal{echo}}\approx 3.5\,\mu\textnormal{s}$ in the lower sweet spot. The flux line bandwidth is measured to be at least $1.3\,\textnormal{GHz}$, potentially limited by a low-pass filter on the line [50], allowing fast transmon frequency adjustment in the nanosecond timescale. This device provides a more compact alternative to the microstrip flux line design shown in recent work [32], allowing more convenient integration of additional circuits such as the readout resonator and cascaded filtering on the same chip.
@@
The configuration of our CPW-based filtering structure is carefully optimised to reduce the energy loss of cavity and transmon through the flux line. We use two 33rd-order low-pass Chebyshev filters [41] in series and based on finite element simulations, they achieve $98\,\textnormal{dB}$ attenuation at the cavity frequency of $\omega_{cav}/2\pi=6.868\,\textnormal{GHz}$ and at least $50\,\textnormal{dB}$ attenuation in the transmon frequency range of $\omega_{T}/2\pi=5.894-7.634\,\textnormal{GHz}$. As a result, energy loss through the flux line is heavily suppressed for both quantum memory and transmon. We verify this by preparing the cavity in a coherent state and tracking the vacuum state $|0\rangle$ population over time at a few different cavity-transmon detunings as shown Fig. 2b. The lifetime reaches a maximum of $T_{1}=590\pm 40\,\mu\textnormal{s}$ at a detuning of $\Delta=\omega_{cav}-\omega_{T}=2\pi\times 260\,\textnormal{MHz}$, at which point the inverse-Purcell decay is suppressed. These photon loss values are comparable to those of typical 3D bosonic superconducting devices without flux integration.
Figure 3: State preparation in the resonant regime. Wigner tomography and a 1D cut along $\mathrm{Re}[\beta]$ of states created using the flux-activated resonant interactions, together with corresponding steps activated to excite the JC ladder. (a) state $\tfrac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$, and (b) state $\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ are created using sequential drives. (c) Fock $|5\rangle$ created from vacuum by simultaneously driving multiple JC transitions.
@@
Our device combines a highly coherent bosonic mode with fast tunability of a nonlinear ancillary element, making it an ideal testbed for resonant control protocols. We first showcase the efficacy of this control framework by preparing arbitrary superpositions of Fock states with a set of analytical pulses. For example, we prepare $\frac{1}{\sqrt{2}}\left(|0\rangle|+|3\rangle\right)$ (Fig. 3a) using a pulse sequence consisting of a $\pi/2$-pulse from $|0\rangle$ to $|1+\rangle$ followed by two $\pi$-pulses to $|2$-$\rangle$ and $|3+\rangle$. The frequencies of each transition are determined using standard spectroscopy experiments, and the pulses durations are chosen to maximize the photon-number parity contrast after climbing each step of the JC ladder [50]. The optimal speed of the adiabatic detuning is calibrated by preparing $|1+\rangle$ on resonance, applying a flux pulse with a smooth, variable-duration ramp until $\Delta=100\,\textnormal{MHz}$, and then measuring the cavity $|1\rangle$ population with a photon-number-resolved spectroscopy. We pick a minimal ramp duration of $200\,\textnormal{ns}$ for which $|1\rangle$ population is maximum. We apply the same method to create $\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ (Fig. 3b), where the phase of the superposition is implemented by simply changing the phase of the pulse addressing this transition. The total pulse sequence to prepare the state in Fig. 3a (b) takes 640 ns (688 ns), and the corresponding fidelity of the final state is estimated as $\textnormal{93}\pm\textnormal{3}\%$ ($\textnormal{89}\pm\textnormal{2}\%$) from density matrix reconstruction [28, 29], limited by transmon decoherence [50].
Figure 4: Givens rotations between two Fock levels. (a) The $|1+\rangle$ state population is swapped with $|3$-$\rangle$ state using a $J_{x}$ pulse, which is then followed by a sideband pulse to $|4+\rangle$ and a second $J_{x}$ pulse that reverts the first swap. This sequence effectively implements an operation between states $|1+\rangle$ and $|4+\rangle$ that translates into a Givens rotation between Fock states $|1\rangle$ and $|4\rangle$ after adiabatic detuning. (b) Measurement of the normalized final state parity for arbitrary rotation angle $\theta$ from an initial state $|1\rangle$, showing the protocol can create arbitrary superpositions of the two Fock states. (c) Wigner plots and Fock states populations obtained from the reconstructed density matrices of the cavity state for $\theta=0,\pi/4,\pi/2$ and $\pi$. The measured population distribution agrees well with the simulated results with real experimental parameters.
@@
As an alternative to sequentially climbing the JC ladder, several transitions can also be activated simultaneously to realize more complex dynamics. For instance, the analog of the spin angular momentum $J_{x}$ over a subspace of dimension $d$ of the JC Hamiltonian can be constructed by driving each transition with amplitude
@@
where A is the overall pulse amplitude and $\eta_{n}$ is an experimentally calibrated scaling factor of the drive strength associated to each transition [50]. Driving $J_{x}$ for a total time of $\pi/A$ enacts pairwise state transfers between target levels. Here we use this multi-tone drive to create Fock $|5\rangle$ from the ground state. We allow $A_{n}$ to assume complex values and obtain the amplitudes using a gradient descent algorithm which maximizes the state preparation fidelity for a fixed time [49, 50]. We present, in Fig. 3c, the measured Wigner function of a Fock $|5\rangle$ prepared with this method using a $800\,\textnormal{ns}$ pulse. We obtain a state fidelity of $\textnormal{75}\pm\textnormal{2}\%$, limited by both the transmon decoherence and imperfection calibrations of $\eta_{n}$ [50].

### v2

@@ Abstract
Universal control of bosonic degrees of freedom provides a hardware-efficient route for quantum information processing with high-dimensional systems. Bosonic circuit quantum electrodynamics (cQED), which leverages auxillary transmons to coherently control long-lived superconducting cavities, is well suited to this goal. However, such systems are traditionally operated in the dispersive regime, where the nearly degenerate cavity transitions prohibit the direct addressability of individual excitation levels of the bosonic mode and increase gate complexity. Here, we achieve direct oscillator control by dynamically accessing the resonant Jaynes–Cummings (JC) interactions, implemented with a hardware that integrates on-chip broadband magnetic flux delivery with a bosonic memory housed in a 3D superconducting cavity with lifetime exceeding 0.5 ms. We demonstrate deterministic preparation of Fock states and their superpositions within 100s of nanoseconds by directly climbing the JC ladder, and realise efficient arbitrary rotations between any pair of Fock states. Our resonant control scheme provides an analytical method for manipulating the entire Hilbert space of the bosonic mode at rates fundamentally more favourable than traditional dispersive strategies. This on-demand access to JC interactions opens a promising path toward realising robust Fock-basis qudits and harnessing the rich dynamics of high-dimensional bosonic systems for quantum information processing.
Effective control over a large Hilbert space is essential for both practical quantum information processing [61, 76, 22] and explorations of the rich underlying concepts of quantum physics [2, 21, 1]. A promising route towards this goal is to leverage the intrinsically large Hilbert space of quantum harmonic oscillators. Notably, bosonic circuit QED systems use 3D superconducting cavities coupled to auxillary transmons to realise highly controllable and coherent oscillators. Such devices have demonstrated Fock state preparation up to 100 photons [11] and cat states with an average photon population of $1024$ [52]. Compared to parallel pursuits using discrete $d$-level systems (qudits) [37, 71], such as molecular spins [58, 54], optical photons [72, 28], trapped ions [42, 60], silicon nuclear spins [75] and superconducting artificial atoms [73, 41, 9], encoding information in higher levels of the oscillator does not introduce additional decoherence channels and is readily compatible with a broad range of existing bosonic quantum error correction schemes [65, 16].
@@
In cQED systems, universal control of the bosonic mode is typically achieved via dispersive coupling to a nonlinear auxillary element, commonly in the form of a transmon qubit [5]. In this regime, oscillator transitions are close in frequency and not individually addressable, as illustrated (Fig. 1a). Therefore, arbitrary operations on the oscillator state often require driving the transmon and the cavity simultaneously with complex numerically optimised waveforms [32, 18, 17]. However, these numerical pulses are susceptible to parasitic dynamics, and their ‘black-box’ nature often complicates effective troubleshooting. Although operations can also be implemented using higher-order parametric processes without numerical optimization [23, 29], they are limited by the achievable coupling strength or drive-induced errors [69, 45]. Therefore, an analytical and versatile control method that directly addresses individual cavity transitions would fill an important gap for bosonic quantum information processing.
@@
In this work, we use flux-activated resonant Jaynes-Cummings (JC) interactions to access number-selective cavity transitions and implement arbitrary yet analytically simple control of a bosonic memory mode. These dynamics are enabled by adjusting the transmon-cavity detuning on demand via a broadband on-chip flux line. This allows us to rapidly switch the coupled system between a highly nonlinear resonant dynamics with well-resolved JC transitions and the dispersive regime with suppressed cavity photon loss. We protect the coherence of the bosonic memory against flux line leakage by incorporating on-chip low-pass filtering over the frequency band occupied by the cavity and the transmon. As a result, the lifetime of the bosonic mode reaches $T_{1}=590\pm 40\,\mu\textnormal{s}$ ($Q\sim 2.6\times 10^{7}$) when the transmon is tuned to be dispersively coupled – more than a two-fold improvement over previous implementations using the magnetic hose [68, 3]. We leverage these hardware capabilities to demonstrate the direct creation of Fock states and their superpositions by sequentially climbing the JC ladder. Furthermore, we demonstrate state transfer between any two chosen energy levels of the cavity using multi-tone pulses and realise arbitrary rotations between them. This resonant control scheme can readily be applied to the general class of spin-boson systems, such as opto-acoustic modes coupled to artificial atoms [4] or motional degrees of freedom of ions controlled by a spin [48]. Thus, our results establish a valuable building block for efficient universal control and quantum information processing using bosonic modes in different physical platforms.
Figure 1: Comparison of the dispersive and resonant regimes. (a) The bosonic transitions in the dispersive regime are degenerate within each qubit manifold and are separated by dispersive coupling strength, $\chi$, between the $|\mathord{\downarrow}\rangle$ and $|\mathord{\uparrow}\rangle$ manifolds. (b) This degeneracy is lifted in the resonant JC regime, with level splittings $\propto g\gg\chi$, enabling fast selective excitation of individual bosonic levels. (c) The frequency spectra of both coupling regimes shown on the same axis.
The resonant control technique explored in this work is based on the theoretical proposal by Strauch [63], which adiabatically connects the eigenstates of the resonant JC Hamiltonian to the product eigenbasis of the decoupled circuits. It improves upon earlier works [38, 20] that rely on sequentially turning on and off the spin-boson coupling.
The dynamics of a spin-boson system on resonance is well approximated within the rotating wave approximation by the JC Hamiltonian [26]
$\hat{H}_{\textnormal{JC}}=g\left(\hat{a}^{\dagger}\hat{\sigma}^{-}+\hat{a}\hat{\sigma}^{+}\right),$ (1)
@@
where $\hat{a}$ ($\hat{a}^{\dagger}$) is the annihilation (creation) operator of the boson in the form of a superconducting cavity, $\hat{\sigma}^{-}$ ($\hat{\sigma}^{+}$) is the lowering (raising) Pauli operator of the spin, realised with a transmon qubit, and $g$ is the coupling strength between them. The $\hat{H}_{\textnormal{JC}}$ eigenstates are the ground state $|0,\,\downarrow\rangle$ and the $N$th-excited levels
@@
$|N\pm\rangle=\frac{1}{\sqrt{2}}\left(|N,\,\downarrow\rangle\pm|N\mathrm{-}1,\,\uparrow\rangle\right),$ (2)
@@
where $|N,s\rangle$ is the product of the cavity Fock state $|N\rangle$ and the transmon state $|s=\,\downarrow$$\uparrow\rangle,$ with eigenfrequencies $\omega_{|N\pm\rangle}=\pm\sqrt{N}g$. This photon-number dependence of $\omega_{|N\pm\rangle}$ leads to the direct addressability of the transitions as illustrated in Fig. 1(b).
@@
$|N+\rangle\,\overset{\omega_{N,r}}{\longleftrightarrow}\,|(N\mathrm{+}1)-\rangle,$ (3)
$|N-\rangle\,\overset{\omega_{N,b}}{\longleftrightarrow}\,|(N\mathrm{+}1)+\rangle,$
with transition frequencies $\omega_{N,r}=-(\sqrt{N\mathrm{+}1}+\sqrt{N})g$ and $\omega_{N,b}=(\sqrt{N\mathrm{+}1}+\sqrt{N})g$ relative to the rotating frame, respectively. Transitions $\omega_{1,b}$, $\omega_{2,r}$, and $\omega_{3,b}$ are shown with orange, green, and teal, in Fig. 1. This set of transitions, together with ground state excitations $|0,\,\downarrow\rangle\leftrightarrow|1\pm\rangle$, can selectively connect any two eigenstates of the resonant JC Hamiltonian with a spectral resolution on the order of $g$ (Fig. 1c). That makes the timescale of our protocol fundamentally more favourable compared to control schemes using the dispersive coupling, as discussed in [64].
@@
To translate this fast control on the JC levels to effective operations on the bosonic mode alone, we simply need to adiabatically decrease the transmon frequency away from the cavity resonance. This transforms the coupled eigenstates in Eq. (2) into products of cavity and transmon eigenstates: $|N+\rangle\to|N,\,\downarrow\rangle$ and $|N-\rangle\to|N\mathrm{-}1,\,\uparrow\rangle$. The red sidebands in Eq. (3) map to the transitions $|N,\,\downarrow\rangle\leftrightarrow|N,\,\uparrow\rangle$, similar to the standard transmon excitations in the dispersive regime [59]. In contrast, the blue sideband, $|N\mathrm{-}1,\,\uparrow\rangle\leftrightarrow|N\mathrm{+}1,\,\downarrow\rangle$, is only directly accessible in the resonant JC dynamics [63]. Therefore, the on-demand access to the resonant JC dynamics, together with the ability to adiabatically detune the transmon from the oscillator, allows us to orchestrate any desired operations on bosonic mode.
@@
To leverage this resonant control scheme, we construct a new bosonic circuit QED hardware that can simultaneously provide strong and broadband flux control while protecting the cavity from decay (Fig. 2a). Our device features a superconducting stub cavity as the long-lived quantum memory coupled to a transmon qubit with a coupling strength $g/2\pi=12.2\pm 0.1\,\textnormal{MHz}$. It is combined with a lithographically defined on-chip flux line placed in close proximity to the SQUID loop with area $470\,\mu\textnormal{m}^{2}$. Flux delivery is highly efficient, featuring a mutual inductance of $0.387\,\Phi_{0}/\textnormal{mA}$. This allows us to minimise the loop area and hence reduce the impact of flux-noise on the transmon coherence times [64]. In this particular device, we measure transmon $T_{1}\approx 23\,\mu\textnormal{s}$ and $T_{2}^{\textnormal{echo}}\approx 3.5\,\mu\textnormal{s}$ in the lower sweet spot. The flux line bandwidth reaches at least $1.3\,\textnormal{GHz}$, allowing fast transmon frequency adjustment in the nanosecond timescale. Although we did not calibrate at higher frequencies due to the presence of a 2 GHz low-pass filter on the coaxial line, our finite-element simulations predict the bandwidth $\sim\textnormal{4}$ GHz, limited by the on-chip low-pass filter [64].
Figure 2: Architecture for robust flux-delivery in a high-Q bosonic circuit QED device. (a) High-Q $\lambda/4$ coaxial cavity coupled to an asymmetric SQUID-transmon (shown in bottom-left inset). The flux bias is delivered by a transmission line wirebonded to an SMA port at the opposite end of the chip. The flux line and a readout resonator partially adopt a coplanar waveguide (CPW) architecture. They are equipped with two 33rd-order Chebyshev low-pass filters in series (schematics shown in top-right inset) and a Purcell filter, respectively, to protect the cavity and the transmon against decay. (b) The cavity $T_{1}$ is measured while the transmon is parked at a few selected flux points with different values of $\Delta=(\omega_{\textnormal{cav}}-\omega_{\textnormal{T}})/2\pi$, showing an increase in $T_{1}$ as the transmon-induced decay is suppressed at larger detunings.
@@
This CPW-based chip design is carefully optimised to reduce the energy loss of cavity and transmon through the flux line. In comparison to the microstrip flux line shown in a recent work [40], our strategy provides a more compact solution that allows convenient integration of additional circuits such as the readout resonator and cascaded filtering on the same chip. We use two 33rd-order low-pass Chebyshev filters [56] with cutoff frequency of 4 GHz in series, schematically depicted in top-right inset of Fig. 1a. Based on finite element simulations, this structure achieves $98\,\textnormal{dB}$ attenuation at the cavity frequency of $\omega_{\textnormal{cav}}/2\pi=6.868\,\textnormal{GHz}$ and at least $50\,\textnormal{dB}$ attenuation in the transmon frequency range of $\omega_{\textnormal{T}}/2\pi=5.894-7.634\,\textnormal{GHz}$. As a result, energy loss through the flux line is heavily suppressed for both quantum memory and transmon. We verify this by preparing the cavity in a coherent state and tracking the vacuum state $|0\rangle$ population over time at a few different cavity-transmon detunings as shown in Fig. 2b. The lifetime reaches a maximum of $T_{1}=590\pm 40\,\mu\textnormal{s}$ at a detuning of $\Delta=(\omega_{\textnormal{cav}}-\omega_{\textnormal{T}})/2\pi=260\,\textnormal{MHz}$, at which point the inverse-Purcell decay is suppressed. These photon loss values are comparable to those of typical 3D bosonic superconducting devices without flux integration.
@@
Our device combines a highly coherent bosonic mode with fast tunability of a nonlinear auxiliary element, making it an ideal testbed for resonant control protocols. We first showcase the efficacy of this control framework by preparing arbitrary superpositions of Fock states with a set of analytical pulses. For example, we prepare $\frac{1}{\sqrt{2}}\left(|0\rangle|+|3\rangle\right)$ (Fig. 3a) using a pulse sequence consisting of a $\pi/2$-pulse from $|0\rangle$ to $|1+\rangle$ followed by two $\pi$-pulses to $|2$-$\rangle$ and $|3+\rangle$. The frequencies of each transition are determined using standard spectroscopy experiments, and the pulses durations are chosen to maximise the photon-number parity contrast after climbing each step of the JC ladder [64]. We calibrate the optimal speed of the adiabatic detuning by first preparing $|1+\rangle$ on resonance, followed by a flux pulse with a smooth, variable-duration ramp until the system is in a strong dispersive regime with $\Delta=100\,\textnormal{MHz}$. We then directly measure the cavity $|1\rangle$ population using the photon-number-resolved spectroscopy and choose a minimal ramp duration of $200\,\textnormal{ns}$ that maximises the population of $|1\rangle$. We further showcase the ease of orchestrating arbitrary multi-component superpositions of Fock states by creating $\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ (Fig. 3b) using the same method. The phase of the superposition is implemented by simply changing the phase of the pulse addressing the corresponding transition. This technique can be directly used to create arbitrary logical states superpositions in the binomial encoding [50, 36]. The total pulse sequence to prepare the state in Fig. 3a (b) takes 640 ns (688 ns), and the corresponding fidelity of the final state is estimated as $\textnormal{93}\pm\textnormal{3}\%$ ($\textnormal{89}\pm\textnormal{2}\%$) from density matrix reconstruction [34, 35], limited by transmon decoherence [64].
@@
As an alternative to sequentially climbing the JC ladder, several transitions can also be activated simultaneously to realise more complex dynamics. For instance, the analog of the spin angular momentum $J_{x}$ over a subspace of dimension $d$ of the JC Hamiltonian can be constructed by driving each transition with amplitude
@@
where A is the overall pulse amplitude and $\eta_{n}$ is an experimentally calibrated scaling factor of the drive strength associated to each transition [64]. Driving $J_{x}$ for a total time of $\pi/A$ enacts pairwise state transfers between target levels. Here we use this multi-tone drive to create Fock state $|5\rangle$ from the ground state. In principle, the optimal amplitudes of each spectral component can be analytically derived to implement a perfect state transfer [63]. In practice, to improve the fidelity beyond the analytical solution, we allow $A_{n}$ to assume complex values and obtain the amplitudes using a gradient descent algorithm which maximises the state preparation fidelity for a fixed time [63, 64]. We present, in Fig. 3c, the measured Wigner function of a Fock $|5\rangle$ prepared with this method using a $800\,\textnormal{ns}$ pulse. We obtain a state fidelity of $\textnormal{75}\pm\textnormal{2}\%$, limited by both the transmon decoherence and imperfect calibrations of $\eta_{n}$ [64].
Figure 3: State preparation in the resonant regime. Left panel: Normalised Wigner plots and 1D cuts along $\mathrm{Re}[\beta]=0$ of states created using the flux-activated resonant interactions. Right panel: corresponding transitions activated to climb the JC ladder. (a) State $\tfrac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$, and (b) state $\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ are created using sequential drives. (c) Fock state $|5\rangle$ created from vacuum by simultaneously driving multiple JC transitions. The progress bars illustrate the steps of state preparation, and the circles indicate the resulting Fock populations.
Figure 4: Givens rotation $\hat{\mathcal{R}}_{1,4}(\theta)$. (a) The $|1+\rangle$ state population is swapped with $|3$-$\rangle$ state using a $J_{x}$ pulse, which is then followed by a sideband pulse to $|4+\rangle$ with an effective rotation angle $\theta$ and a second $J_{x}$ pulse that reverts the first swap. This sequence effectively implements a gate between states $|1+\rangle$ and $|4+\rangle$ that translates into a Givens rotation between Fock states $|1\rangle$ and $|4\rangle$ after adiabatic detuning. (b) Normalised final state parity measured for arbitrary rotation angle $\theta$, showing that the protocol can create arbitrary superpositions of two Fock states. (c) Normalised Wigner function plots and Fock states populations of the cavity state for $\theta=0,\pi/4,\pi/2$, and $\pi$. The population distributions are obtained from the reconstructed density matrices and consistent with both the ideal state (target) and the outcomes simulated with real experimental parameters (black).